### v1

# Learning to Synthesize Fault-Tolerant Quantum Circuits by Diagonalization

Abstract
Compilation of quantum programs into circuits expressed with discrete gate sets is essential for fault-tolerant quantum computing. Optimal methods for discovering high-precision implementations of unitaries in discrete gate sets such as the Clifford+T gate set are intractable. Search-based synthesis methods, including Reinforcement Learning (RL) and simulated annealing, are promising as they empirically discover efficient implementations of low-depth unitaries. We leverage search-based methods to reduce the general unitary synthesis problem to one of synthesizing diagonal unitaries; a problem solvable efficiently in general and optimally in the single-qubit case. Relying on several improvements in deep learning architectures essential for quantum datasets, our RL-based approach demonstrates how search-based methods and mathematical decompositions can be combined to efficiently find high precision implementations of unitaries taken from an array of real quantum algorithms. On these benchmarks we observe up to an average of $94\%$ fewer non-Clifford gates compared to the Quantum Shannon Decomposition.
Index Terms:
quantum compilation, reinforcement learning, unitary synthesis, interpretability
I Introduction
Recent small scale demonstrations of logical qubits [1, 2] bode well for the future of Fault Tolerant (FT) quantum computing. In this regime, circuits are expressed in a discrete gate set, usually a small set of Clifford gates (e.g. H, S, CZ) augmented with a non-Clifford gate (e.g. T or V) for universality. The Clifford+T gate set can approximate any unitary transformation [3] and it has been used as the target of many FT quantum compilation algorithms [4, 5, 6, 7, 8]. As executing non-Clifford gates in many FT frameworks requires expensive pre-processing in the form of magic state distillation [9, 10, 11], a common goal of code generation tools is minimizing the number of non-Clifford gates in the circuit. Besides circuit quality (i.e. gate count), the scalability of FT compilers with respect to circuit size and required precision are crucial.
It is known that if each entry of a unitary are elements of the ring $\mathbb{Z}[i,\frac{1}{\sqrt{2}}]$, that unitary can be implemented exactly with the Clifford+T gate set [12, 13]. Unitaries which appear in real algorithms are rarely this well structured, so they must be approximately synthesized instead. Analytical decomposition [14, 15] methods for Clifford+T generation are universally applicable, but circuit quality may suffer at high precision. In restricted cases, optimal approximate implementations of unitaries in discrete gate sets can be efficiently found [16]. Beyond single-qubit rotations restricted to a single axis, finding optimal approximate implementations of unitaries in the Clifford+T gate set is intractable [17].
Search-based circuit synthesis methods perform discrete searches through the space of possible quantum gate placements [18, 19, 20]. Since this problem can be expressed as a Markov Decision Process, Reinforcement learning (RL) is a natural fit [21, 22, 23, 24, 25]. Search based algorithms can discover more resource efficient circuits than analytical methods, but fail to converge when compiling high precision implementations of complex unitaries. This is because continuously parameterized single-qubit rotation operations typically correspond to long sequences of discrete gates (see Figure 1). To ensure high circuit quality, these algorithms should be used in practice along with analytical methods.
We provide two contributions which enable hybrid search- and analytical-based methods to compile complex unitaries:
  1. 1.
    We bypass the need to transpile difficult rotations with search-based tools by loosening the synthesis objective from matrix inversion to diagonalization. This allows search-based methods to synthesize $n$-qubit unitaries up to $2^{n}$-$1$ continuous degrees of freedom. Analytical decompositions and optimal single-qubit techniques are used to implement complex rotations. Diagonalization produces high precision results where other search-based tools fail. This process is general; other search-based tools can be retrofitted for diagonalization.
  2. 2.
    We introduce modifications to standard neural network architectures that enable machine learning methods to better tackle problems involving quantum datasets. Our Deep Value Network for Unitaries (DVNU) approach advances the state-of-the-art for RL based synthesis methods by improving the complexity of circuits which can be synthesized using deep learning.
Our data indicate that the relaxation to diagonalization along with our neural architectural changes enable efficient and high precision synthesis of unitaries from real quantum algorithms. Our approach uses RL as its fast inference time makes it practical for use in full stack quantum compilers.
Figure 1: Strategies for synthesizing unitaries into Clifford+T gates. Analytical decomposition approaches, such as the Quantum Shannon Decomposition are universally applicable but often produces poor results. Search-based methods are able to find very high quality (low gate count) approximations, but are intractable except at low precision. Diagonalization is a hybrid approach that enables high quality and high precision. gridsynth [16] is used to optimally decompose continuous $R_{Z}$ operations (colored boxes) into Clifford+T gates.
The remainder of the paper is organized as follows: Section II discusses necessary background information relating to synthesis with discrete gate sets. Section III explains how matrix diagonalization can be framed in the language of synthesis. Section IV introduces neural network architectural developments natural for quantum datasets. Section V evaluates the new architecture on synthesis tasks and shows how our RL based approach can effectively synthesize many unitaries taken from partitioned quantum algorithms. Section VI discusses methods of scaling, interprets the learned value function, and argues for the utility of RL in unitary synthesis. Finally, section VII offers concluding remarks.
II Background
II-A Discrete and Fault-Tolerant Gate Sets
Universal quantum computers must represent complex unitaries on $n$ qubits with high fidelity. Modern noisy machines offer universal basis gate sets, which compilers use to express unitary operators. The fidelity with which these machines express algorithms is highly dependent on the choice of gate set [26]. Universal gate sets typically consists of an entangling two-qubit gate (e.g. CNOT or CZ gates) along with a set of fixed and continuously parameterized single-qubit gates (e.g. $X$ and $R_{Z}(\theta)$).
Quantum Error Correcting Codes encode the state of one or more logical qubits into the states of many physical qubits [27]. Some logical operations require only that the same physical operation be applied to each physical qubit. These transversal gates form closed finite groups and are not universal [28]. The Clifford group coincides with the transversal gate set of many codes of interest. For universal quantum computation, the logical gate set must contain a non-transversal operation. The T gate is often considered as, being a member of the third level of the Clifford Hierarchy, it can be executed through magic state injection and Clifford gate corrections [29]. This work focuses on approximately compiling unitaries into the Clifford+T gate set.
Continuous rotations such as the $R_{Z}(\theta)$ gate are often unavoidable in quantum algorithms. For some precision $\epsilon$, translating these operations into Clifford+T gates requires $O(\log\frac{1}{\epsilon})$ gates. Except in select cases, this is true regardless of the angle of rotation. $R_{Z}$ rotations can be approximated optimally to arbitrary precision using the gridsynth algorithm [16]. Other methods also exist to handle these rotations [30, 31, 32], but for the scope of this paper we consider only the optimal ancilla free case.
II-B Circuit Synthesis
Say that we want to synthesize the unitary labeled “Target” in Figure 1. Any unitary can be analytically decomposed into a sequence of CNOT and general single qubit unitary gates. The Quantum Shannon Decomposition and its variants are able to produce circuit implementations which approach the asymptotic lower bound of $O(4^{n})$ CNOT and $R_{Z}$ gates [14, 15]. Although these analytical approaches can always be applied and allow for arbitrarily high precision, they often produce circuits which contain many non-Clifford gates (i.e. $R_{Z}$ rotations). These $R_{Z}$ rotations can be implemented in several ways, including optimal decomposition into Clifford+T gates using gridsynth.
Other algorithms often adopt a search-based approach. Here, a guiding function is used to iteratively modify a circuit, often gate-by-gate, until the target unitary is implemented. Examples of such algorithms include Reinforcement Learning approaches [21, 22, 23, 24, 25]. simulated annealing [20], and several optimal and heuristic synthesis algorithms [18, 17, 19]. Empirically, these state-of-the-art methods are able to find very efficient implementations of unitaries. Search-based tools are restricted in that they only operate with discrete gate sets, and therefore have no direct way of handling continuous rotations such as $R_{Z}$ gates. For unitaries which contain these rotations, search-based tools are only able to find low precision implementations.
Analytical and search-based approaches both seek to directly invert the adjoint of the target unitary. We propose an alternative approach: diagonalization (Figure 1). The diagonalization process consists of a two-headed search which halts when the target’s adjoint is diagonalized (i.e. not fully inverted). Diagonalizing ensures that up to $12^{n}-$ 1 continuous $R_{Z}$ operations can be handled. Although not optimal, this approach is able to find high precision implementations of complex unitaries (unlike search-based inversion methods), and produces significantly higher quality outputs than pure analytical methods (see Table II). In the worst case, diagonalization acts exactly like the underlying search-based algorithm which implements it.
II-C Formalizing Search-Based Quantum Circuit Synthesis
The goal of a quantum circuit synthesis algorithm is to find a quantum circuit implementation of a target unitary matrix. This problem is typically framed as finding a circuit which inverts the adjoint of the target.
When solving the synthesis problem directly by matrix inversion, we assume that every circuit can be described as a sequence of primitive gates taken from a finite set $A\subset\mathbb{U}(2^{n})$. Every primitive gate has a known associated unitary $a\in A$. A circuit consisting of $t$ gates can therefore be associated with a unitary matrix
$C_{t}=a_{t}\times\dots\times a_{1}.$ (1)
A quantum circuit with associated unitary $C_{t}\in\mathbb{U}(2^{n})$ is said to implement a target unitary matrix $U_{tar}\in\mathbb{U}(2^{n})$ when the distance condition
$d_{HS}(C_{t},U_{tar})=\sqrt{1-\frac{1}{4^{n}}|Tr(C_{t}U_{tar}^{\dagger})|^{2}}\leq\epsilon$ (2)
is satisfied. Here $\epsilon\ll 1$ is a hyperparameter which determines how precisely a target must be implemented. The distance condition in Equation 2 is invariant to global phase. This is a desirable property as quantum processes that differ only by a global phase cannot be distinguished by projective measurements [29].
II-D Synthesis as a Markov Decision Process
Figure 2: The agent observes the state of the environment $s_{t}$, uses the policy to commit to an action $a_{t}$, then receives a reward $r_{t}$. Model-based synthesis approaches can evaluate the value of many actions before committing to them.
Markov Decision Processes (MDPs) provide a framework for describing stochastic decision problems. Posing quantum circuit synthesis as an MDP has been demonstrated numerous times in the literature [21, 22, 23, 24, 25] We define the MDP as a tuple $(S,S_{I},S_{T},A,r,p)$. Here $S$ is a set of states, $S_{I}$ a set of initial states, $S_{T}$ a set of terminal states, $A$ a set of actions or gates, $r:S\times A\rightarrow\mathbb{R}$ a reward function, and $p$ a distribution which assigns a probability to next states given a state-action pair.
Given an MDP, a Reinforcement Learning (RL) algorithm optimizes for a policy which maximizes the total expected cumulative rewards. Formally, a policy denoted as $\pi(a_{t}|s_{t})$ assigns a probability to each possible action $a_{t}\in A$ given state $s_{t}$. A state $s_{t}$ has value
$V^{\pi}(s_{t})=\mathbb{E}_{\pi}\sum_{k=0}^{T}\gamma^{k}r(s_{t+k},a_{t+k})$ (3)
where $\gamma\in(0,1]$ is a discount factor. The value is the expected cumulative discounted reward if policy $\pi$ is followed. We can additionally define the state action value or Q function
$Q^{\pi}(s_{t},a_{t})=r(s_{t},a_{t})+\gamma\mathbb{E}_{p}V^{\pi}(s_{t+1})$ (4)
which is the value of taking action $a_{t}$ when in state $s_{t}$ and then following policy $\pi$.
RL algorithms typically consist of two entities, an agent and an environment (illustrated in Figure 2). The agent is trained in order to maximize the cumulative rewards it receives from the environment. Learning the optimal value function leads to a synthesis policy which produces circuit implementations of unitaries that contain the fewest possible gates. For the scope of this paper, it is assumed that a neural network is used to learn either a value function (Equation 3) or a Q function (Equation 4). A policy can be obtained by simply taking the action which maximizes either function.
Table I compares various unitary synthesis approaches which can be used to target the Clifford+T gate set. Among search-based methods, the simulated annealing based Synthetiq tool empirically finds better implementations of unitaries than other methods [20]. Even so, Synthetiq fails to produce results for high precision implementations of complex unitaries as the space of circuits is too large to search. Synthesizing complex unitaries to high precision requires a mechanism for handling continuous rotations; this is possible in both analytical and diagonalization approaches.
Method Qubits Category Precision Non-random?
MCTS [25] 2-5 Search (RL) $10^{-1}$ No
Policy Iter. [24] 1 Search (RL) $10^{-2}$ No
DQN [23] 1-2 Search (RL) $10^{-3}$ No
Synthetiq [20] 1-4 Search (SA) $10^{-3}$ Yes
DVNU (ours) 1-3 Search (RL) $\leq 10^{-6}$ Diagonalizable
gridsynth [16] 1 Analytical - $R_{Z}$
QSD [14] 2+ Analytical - Yes
Table I: Unitary synthesis approaches for the Clifford+T gate set. Among RL approaches, DVNU handles more complex inputs and with diagonalization can synthesize to very high precision. Precision is Hilbert-Schmidt distance. Non-random indicates that unitaries from real benchmarks can be handled. Synthetiq uses Simulated Annealing and empirically outperforms other pure search-based tools at low precision. gridsynth optimally decomposes 1-qubit $R_{Z}$ unitaries to any precision. The QSD is an analytical method which can be used along with gridsynth to synthesize Clifford+T circuits.
III Diagonalization
Directly synthesizing a unitary by finding a sequence of gates which inverts it is intractable when high-precision is needed. This is because a large number of discrete gates are required to implement continuous rotations. To address this we shift the problem scope from inversion to diagonalization. This approach assumes there are continuous degrees of freedom in unitaries which cannot be handled by a search-based synthesis process. To fit the paradigm of search based synthesis as described in Section II-C, we first define the problem of diagonalization as an MDP. At time $t\in\{1,\dots,T\}$ in the synthesis process, the state $s_{t}\in S$ is defined by
$s_{t}=L_{t}U_{tar}^{\dagger}R_{t}$ (5)
where $L_{t},R_{t}$ are each sequences of discrete operations (as in Equation 1). The set of initial states contains the adjoint of target unitaries $S_{I}=\{U_{tar}^{\dagger}\}$. The set of terminal states is the set of all states which can be approximately inverted by a diagonal unitary matrix. This corresponds to the set of states satisfying
$d_{D}(s_{t})=\min_{\theta\in\mathbb{R}^{2^{n}}}d_{HS}\big{(}s_{t},\exp(-i\theta\cdot\hat{\sigma})\big{)}\leq\epsilon$ (6)
where $\hat{\sigma}$ is the vector of all $n$-fold tensor products of the Pauli $\sigma_{I}$ and $\sigma_{Z}$ matrices. For brevity’s sake, we denote the diagonal operator resulting from the matrix exponential as $D_{\theta}$. The reward function is an indicator function defined as
$r(s_{t})=\begin{cases}+1&\text{if }s_{t}\in S_{T}\\ -1&\text{else}\end{cases}$ (7)
Given a terminal state $s_{T}$, the corresponding circuit is
$R_{T}s_{T}^{\dagger}L_{T}=R_{T}D_{\theta}^{-1}L_{T}.$ (8)
Figure 3 illustrates the general form of three-qubit circuits which satisfy this structure. As an example, the Toffoli gate can be diagonalized in just two Hadamard gates.
Figure 3: Diagonalization ansatz for three-qubit unitaries. The diagonal matrix corresponds to $12^{n}-$ 1 continuous degrees of freedom which can be efficiently implemented with $R_{Z}(\theta)$ and CNOT gates. The $L$ and $R$ subcircuits are made of discrete Clifford+T operations. A Toffoli only requires two H gates to be diagonalized and its diagonalization can be implemented with only CNOT and $R_{Z}(\pm\frac{\pi}{4})$ ($T$ or $T^{\dagger}$ gates).
This process can be be thought of as a Singular Value Decomposition where the left and right singular vector matrices are restricted to unitaries which can be efficiently implemented by searching through gate sequences from a discrete gate set (i.e. Clifford+T). Although our implementation of synthesis by diagonalization uses RL, other search-based synthesis methods are equally valid.
III-A Handling Diagonal Unitaries and $R_{Z}$ Gates
The diagonal operator $D_{\theta}$ captures continuous degrees of freedom that cannot be efficiently handled by the $L_{t}$ and $R_{t}$ unitaries. An ansatz implementing $D_{\theta}$ can be prepared using techniques from the synthesis diagonal unitaries and of Pauli strings [33, 34, 35]. These diagonal unitaries require only $CNOT$ and $R_{Z}$ gates. We assume that the $R_{Z}$ rotations will be implemented without ancilla in the Clifford+T gate set using gridsynth [16].
IV Reinforcement Learning with Unitaries
To diagonalize unitaries, we adopt a Reinforcement Learning approach called the Deep Value Network for Unitaries (DVNU). DVNU indirectly learns a diagonalization policy $\pi$ using Value Iteration [36]. Compared to the approaches described in [22, 23], our model-based method trains faster and achieves better synthesis performance.
We introduce advancements necessary for compiling more complex quantum programs. First, we introduce a global phase invariance operation. Invariance to global phase ensures that networks learn to associate values with a canonical form of a unitary state. Second, we introduce a unitary perturbation data augmentation. By applying small unitary disturbances to states during training, the network learns to associate values with regions centered around points on the Bloch sphere. The impact of these architectural changes are illustrated in Figure 5. An ablation study summarizes the effect on performance of all architectural changes (see Figure 7).
IV-A Modeling Synthesis
Our search-based synthesis method relies on a form of RL called model-based learning.
Definition 1.
A model is a function $f$ which predicts the dynamics of a controllable environment $f(s_{t},a_{t})=\tilde{s}_{t+1}$.
In the case of quantum circuit synthesis, knowledge of the state $s_{t}$ and the action $a_{t}$ allows for exact, deterministic, knowledge of the system dynamics because $s_{t+1}=a_{t}\times s_{t}$ or $s_{t+1}=s_{t}\times a_{t}$.
IV-B A Global Phase Invariant Network Architecture
Two observables which differ only by a global phase cannot be distinguished experimentally [29]. When dealing with unitary operators numerically, we are forced to choose representations which may differ due to global phase. To ensure that equivalent unitaries are represented in the same way, we require global phase invariance.
Definition 2.
A function $g$ is said to be global phase invariant if for two unitaries $U,W\in\mathbb{U}(2^{n})$ such that $U=e^{i\theta}W$, it is the case that $g(U)=g(W)$.
We construct a new phase invariant neural network layer similar to the manner of [37], which showed how nonlinear activation functions may be made equivariant to complex phase. Figure 4 illustrates this process and a proof of its correctness can be found in the Appendix.
[…]
Figure 4: A unitary $U$ can be transformed to global phase invariant representation $W$ by 1)1)1 ) squaring $U$ element-wise, 2)2)2 ) computing and conjugating the normalized mean to obtain a unit magnitude complex number $\hat{\mu}^{*}$, 3)3)3 ) multiplying $\hat{\mu}^{*}\times U$, and 4)4)4 ) conditionally correcting the phase by an angle of $\pi$.
IV-C Unitary Perturbations for Robust Value Learning
Unitary synthesis constructs approximate implementations of unitaries. As indicated by the distance condition in Equation 2, each unitary is associated with an $\epsilon$-disk of unitaries which are considered to be approximately equivalent. This fact naturally inspires a perturbative data augmentation technique: small unitary disturbances are applied to unitary states in order to learn this approximate equivalence structure.
Concretely, given a unitary state $s_{t}$, a unitary perturbation $U_{\epsilon}$ satisfying $d_{HS}(U_{\epsilon},I)<\epsilon$ is applied so that the state viewed by the agent is $U_{\epsilon}\times s_{t}$ (or $s_{t}\times U_{\epsilon}$). A visualization of the unitary perturbation data augmentation is illustrated in Figure 5. The impact on performance is depicted in Figure 7.
IV-D Neural Radiance Field Encoding
Unitary states are represented using $2\times 2^{n}\times 2^{n}$ real tensors, where $n$ is the number of qubits. As we focus only on cases of high-precision synthesis where $n\in\{1,2,3\}$, our input data is extremely information dense. Past work has shown that neural networks struggle to fit high frequency patterns in data [38, 39]. We adopt a positional encoding technique used widely in the field of novel view synthesis and introduced by the NeRF project [40]. Input values are transformed by a map $\gamma:\mathbb{R}\rightarrow\mathbb{R}^{2L}$ such that
$\begin{split}\gamma(x_{i})=\Big{(}\sin(2^{0}\pi x_{i}),\cos(2^{0}\pi x_{i}),\dots,\\ \sin(2^{L-1}\pi x_{i}),\cos(2^{L-1}\pi x_{i})\Big{)}\end{split}$ (9)
This encoding map plays a vital role in allowing networks to detect small magnitude differences in each element of a unitary state. The encoding therefore transforms input tensors from a shape of $(2,2^{n},2^{n})$ to a shape of $(2,2^{n},2^{n},2L)$. For our experiments, we set $L=15$.
IV-E Deep Value Network for Unitary Synthesis
Figure 5: Visualization of global phase invariance and unitary perturbative data augmentations. Left: Without global phase invariance, the Bloch sphere is rotated so that equivalent states have different representations. Center: With no unitary perturbations on the inputs, reachable states are only represented as single points. Right: Global phase invariance and unitary perturbations ensure that states are represented consistently and that the network learns to associate entire $\epsilon$-regions with values.
We leverage deep value iteration to learn gate set specific value functions. We refer to this architecture as the Deep Value Network for Unitaries (DVNU). This contrasts with past work such as [22, 23] which leverages a Deep Q Network (DQN) [41]. To train a DQN agent, a state-action value (Equation 4) must be learned for each $(s_{t},a_{t})$ pair. To ensure adequate training, the agent must receive examples of many different actions taken in each state. When the number of actions (in this case the gate set $|A|$) is large, the sample inefficiency of the DQN algorithm limits training progress. In comparison, the DVNU must only learn the value of a state $s_{t}$. The policy used is
$\pi(a|s_{t})=\arg\max_{a\in A}V^{\pi}_{\theta}(a\times s_{t})$ (10)
Value iteration [42] is possible because synthesis can be modeled exactly and the reward is a deterministic function of the state. The previous work of [21] also employed a method based on value iteration for topological quantum computing. More recently [24] demonstrated how in restricted cases tabular (as opposed to deep RL) policy iteration can be used to solve for optimal synthesis policies.
The DVNU additionally leverages global phase invariance (Sec IV-B) and unitary perturbative data augmentations during training (Sec IV-C). As with many of these methods, DVNUs can be extended to be topology aware: the agent will only propose gates that are allowed by a machine’s qubit connectivity. As the learned value function is gate set and topology dependent, a new agent must be trained in order to handle new topologies and new basis gates.
Figure 6: Accuracy of global phase invariant and non-invariant architectures as a function of phase shift. Two otherwise identical networks are trained to match unitaries with their associated Clifford+T gates. Global phase shifts (angles on the unit circle) are applied to unitaries during training and testing. The invariant architecture maintains perfect accuracy (keeps a radius of $100\%$); it correctly matches unitaries with their associated gate no matter the global phase.
V Experiments
In this section we evaluate our Reinforcement Learning diagonalization approach to quantum circuit synthesis detailed in Section IV. We show how our method can synthesize high-precision unitaries which are out of reach for other synthesis approaches.
Figure 7: Ablation study for 1-qubit (left) and 3-qubit (right) synthesis. Lines are the mean of four runs, shaded regions depict extrema. The $y$-axis combines agent success and problem difficulty. Global phase invariance can be learned in the simplest cases, but is necessary for harder problems. Invariant architectures outperform non-invariant ones given enough training time.
V-A How does global phase invariance impact performance?
To answer this question, we can compare the ability of invariant and non-invariant architectures to learn useful properties about unitaries. We take as a proxy task the problem of predicting which unitary is associated with which gate in a finite gate set. Two structurally identical multi-layer perceptron [43] networks are trained to assign gate labels to unitaries associated with the two-qubit Clifford+T gate set for 250 epochs. Figure 6 illustrates how the network which first applies the invariance operation detailed in Section IV-B is resilient to global phase disturbances; no matter the global phase shift it is able to correctly match unitaries with their associated gates.
Global-phase invariance enables neural networks which operate on quantum datasets to achieve better performance. Actively addressing this symmetry removes the requirement that networks learn invariant representations. In the 1-qubit case, the impact of global-phase invariance is less evident; agents trained with and without dephasing achieve similar performance. The effect is more pronounced when tackling 3-qubit circuit synthesis (see Figure 7).
V-B How does DVNU compare to other RL approaches?
Figure 8: Probability that a DQN agent (left) and a DVNU agent (right) successfully finds implementations of random 1-qubit Clifford+T unitaries by inversion. The Target Depth-axis shows the number of gates in 250 randomly generated compilation targets. Adding gates increases compilation difficulty. The Iteration Number-axis measures how much training time has passed. The vertical axis shows the probability that a compilation agent will successfully compile a target circuit. Our DVNU agent significantly outperforms state of the art deep learning based unitary synthesis work demonstrated in the literature.
For this experiment, random unitary targets are obtained by randomly generating quantum circuits in the Clifford+T gate set consisting of $\{H,X,Y,Z,S,T,S^{\dagger},T^{\dagger}\}$ gates. Circuits are generated by randomly selecting a member of the 1-qubit Clifford group followed by either $T$ or $T^{\dagger}$. Random circuits are used during training because they can easily be generated and because the difficultly of synthesizing the unitaries associated with these circuits can be tuned by adding or removing gates. Agents in this experiment are trained to invert unitaries, meaning that the state representation used is $s_{t}=C_{t}U_{tar}^{\dagger}$.
The metric of interest here is the likelihood that a discrete gate synthesis algorithm finds a circuit implementation of a target unitary. There are $O(|A|^{d})$ depth $d$ possible circuits to check, so performance on this task estimates how well an agent has learned to navigate this huge state space.
We train two models, a DQN as in [22] and a DVNU. Both networks were made invariant to global phase. Agents are trained for 150 iterations (about 6 hours) on a machine with an AMD EPYC 7763 CPU and an NVidia A100 GPU. Figure 8 illustrates the results of this experiment. After training, the 1-qubit DVNU agent was able to synthesize random Clifford+T circuits of depth 80 with a success rate of 54%. Comparisons to other RL synthesis projects are highlighted in Table I.
Given the same time budget, the DVNU agent achieves higher synthesis performance across all evaluated problem difficulties. This can partially be explained by the greater sample efficiency of value iteration compared to Q learning. By leveraging a deterministic and exact model of synthesis, values are associated with states instead of state-action pairs.
V-C How well can the RL agent diagonalize unitaries?
2 Qubits Heisenberg HHL Hubbard QAOA QPE Shor TFIM VQE Mean
Time per 0.98 0.96 0.97 0.98 0.97 0.96 0.98 0.96 0.97
Unitary (s) 0.93 0.63 0.92 0.86 0.59 0.73 0.52 0.87 0.76
Success Rate 100% 100% 100% 100% 100% 100% 100% 100% 100%
93.7% 37.4% 42.8% 27.1% 29.1% 41.7% 51.0% 20.0% 42.9%
$R_{Z}$ Count 7.66 8.28 7.78 9.08 6.92 7.50 0.14 4.82 6.52
1.0 1.64 0.85 1.06 0.91 1.56 1.0 0.94 1.12
T Count 0.60 0.69 0.09 0.93 1.00 0.33 0.00 0.07 0.46
0.03 0.36 0.00 0.00 1.15 0.10 0.00 0.16 0.23
Clifford Count 17.29 18.06 16.90 18.48 15.47 16.09 8.08 13.32 15.46
4.39 4.01 3.16 2.10 3.50 4.07 5.47 8.31 4.38
3 Qubits Time per 1.07 1.06 1.06 1.18 1.07 1.18 1.07 1.05 1.09
Unitary (s) 1.03 1.07 1.12 1.02 1.06 1.12 1.01 0.98 1.05
Success Rate 100% 100% 100% 100% 100% 100% 100% 100% 100%
27.4% 9.6% 55.6% 11.2% 78.5% 18.0% 9.4% 13.3% 27.9%
$R_{Z}$ Count 44.08 29.91 34.69 30.65 37.54 37.73 41.62 37.97 36.77
1.88 2.06 0.87 2.27 1.48 4.21 1.90 0.72 1.92
T Count 3.02 1.76 3.93 2.11 2.91 2.60 2.62 4.48 2.93
0.10 2.15 0.47 0.00 3.54 0.28 0.17 0.43 0.89
Clifford Count 91.42 72.27 85.92 70.90 87.81 83.40 79.03 98.62 83.67
13.06 15.94 14.85 10.22 14.00 11.92 16.76 16.69 14.18
Table II: Synthesis of two- and three- qubit unitaries taken from partitioned circuits. Compared to the QSD (top rows), the diagonalizing agent (bottom rows) on average reduces the number of $R_{Z}$ gates by 82.8% and 94.8% for two- and three- qubit unitaries respectively. Comparisons are made to the QSD because other synthesis tools fail to find solutions given $\epsilon=10^{-6}$.
Given a 2-3 qubit target, we wish to determine how well the DVNU based Reinforcement Learning diagonalization approach can synthesize unitaries. Throughout these experiments, we require that unitaries be synthesized to a distance of $\epsilon=10^{-6}$ (see Equation 2). At this level of precision, pure search-based tools are unable to synthesize unitaries by inversion. To test this assertion, we attempted to synthesize our suite of unitaries with Synthetiq: a simulated annealing tool which empirically achieves state-of-the-art performance [20]. We relaxed the requirement that Synthetiq fully invert matrices by using the diagonal distance function (Eq 6). Because Synthetiq attempts to invert matrices (i.e. uses a state representation in the form of $s_{t}=C_{t}U_{tar}^{\dagger}$ as opposed to $s_{t}=L_{t}U_{tar}^{\dagger}R_{t}$), many targets remain out of reach. Solutions could only be found within a 20 minute time out period when inputs were already diagonal.
As a point of reference, we compare synthesis results of the diagonalizing agent to the Quantum Shannon Decomposition (QSD) [14]. This analytical decomposition was chosen because no other search-based agent that we tested was able find solutions by matrix inversion with sufficient precision. Results from the QSD are optimized by converting and combining $R_{Z}$ gates into Clifford+T gates when possible.
To train the DVNU diagonalizing agent, we randomly generated separate $A$ and $B$ Clifford+T circuits, then inserted random $R_{Z}$ rotations to form targets in the form $AR_{Z}^{\otimes n}(\theta_{i})B$. In the 2- (3-)qubit case, the $A$ and $B$ circuits contained up to a combined 30 (20) gates.
We evaluated the diagonalizing agent’s performance on a set of unitaries taken from partitioned quantum algorithms. This suite of algorithms includes Shor’s Algorithm [44], TFIM, Heisenberg, and Hubbard model quantum chemistry simulation circuits [45], trained VQE [46] and QAOA [47] circuits, and Quantum Phase Estimation circuits [48]. The VQE and QAOA circuits were generated by MQTBench [49]. Each set of unitaries was filtered to ensure that every unitary was unique. There were 22,323 different 2-qubit unitaries and 45,202 different 3-qubit unitaries. Table II summarizes.
We find that in the two-qubit case, diagonalization can find circuits implementing 42.9% of unitaries across all benchmarks. About 79% of these unitaries are already diagonal, therefore any synthesis agent using a diagonal distance (Equation 6) should find solutions in these cases. Most other unitaries contain 2-8 gates, typically Hadamard gates. These unitaries are often just controlled rotations about a non-Z axis.
For the three-qubit case, realistic unitaries are far more complex: they are more likely to contain unitaries which do not conform to the diagonalization ansatz. Approximately 27.9% of three-qubit unitaries from our suite of partitioned circuits could be synthesized by the diagonalizing agent. Of these, approximately 57% were already diagonal. On average, 3.84 CNOT gates are needed to implement the diagonal operator, meaning the diagonalizing DVNU agent typically finds diagonalizing $L_{t}(\cdot)R_{t}$ implementations with an average of about 11 gates. Some common unitaries are almost trivial to synthesize under the diagonalizing approach. Toffoli (which is synthesized with optimal ancilla-free T-count and illustrated in Figure 3) and $CCR(\theta)$ gates are examples of such unitaries.
Compared to the QSD decomposition, the diagonalizing agent produces solutions with far fewer non-trivial rotation gates. Implementing an $R_{Z}$ gate requires $O(\log\frac{1}{\epsilon})$ T gates when synthesized optimally with gridsynth [16]. For $\epsilon=10^{-6}$ this is approximately 60 T gates per $R_{Z}$. Lone T gates are almost negligible compared to continuous rotations in the high precision regime. Compared to the QSD, the DVNU synthesizer reduces the average number of non-Clifford gates by $82.8\%$ for two qubit unitaries and $94.8\%$ for three qubit unitaries.
Although the success rate for the DVNU agent is lower than for mathematical decompositions, the potential savings in non-Clifford gates and the speed with the process runs remain strong arguments for its utility.
VI Discussion
Section V demonstrates how our model-based circuit synthesis agent performs. In this section, we discuss the implications of these results and pose future research directions.
(a) State transition diagram under a noisy and flat value function.
(b) State transition diagram under the learned value function.
(c) Proximity to the identity ($1-d_{HS}(s_{t},I)$) vs. gate count of random Clifford+T circuits. This value function is noisy and flat, making agents act randomly in most of the state space.
(d) Learned value of synthesis states vs. gate count of random single qubit Clifford+T circuits. This value function guides agents towards terminal states.
Figure 9: The impact of the value function on state transitions. (a) Agents which use noisy and flat value functions to plan actions will act randomly. Nodes $A$ and $B$ represent the initial and terminal states respectively. Darker contours represent increasing value magnitudes. (b) Agents which use learned values to plan actions tend towards terminal states. (c) shows the Hilbert-Schmidt distance which is a noisy and flat value function. Lines are the mean distances of 250 random circuits, the shaded regions are interquartile ranges. In synthesis with discrete gate sets, unitary distance does not correlates with gate count. States which require fewer gates to reach a terminal state generally do not correspond to lower distances than states which require more gates. (d) shows the learned value function which is based on path distance. States which require fewer gates to reach a terminal state have higher values than those which require more gates. Gate cancellation in randomly generated circuits contributes to variance in state values.
VI-A Trade-Offs for High Precision Synthesis of Realistic Unitaries
Unitary synthesis has proven to be an invaluable tool for transpiling near-term quantum circuits [50]. Synthesis algorithms targeting Fault Tolerant gate sets are certain to fit a similar niche. In order to be useful for transpilation, FT synthesis algorithms must be able to approximate unitaries to high levels of precision. In this paper, we have demonstrated how our Reinforcement Learning based diagonalization approach can synthesize realistic unitaries to a precision of $\epsilon=10^{-6}$. This value was chosen as it is 1000×1000$\times$1000 × more precise than previous methods (see Table I). Because the state representation is much more complex in the diagonalization approach ($s_{t}=L_{t}U_{tar}^{\dagger}R_{t}$ instead of $s_{t}=C_{t}U_{tar}^{\dagger}$ ), the search process is much more difficult. To this end, the modifications made to the neural architecture of DVNU (Section IV) are essential for machine learning approaches to perform well. Further innovations may further improve the applicability of deep learning to datasets of unitary matrices.
Because much of the difficulty, and the need for approximation, comes from translating $R_{Z}$ rotations to Clifford+T gates, our diagonalizing tool can produce implementations of many unitaries even more precise than this. Targeting higher levels of precision using DVNU often requires retraining, which takes approximately 2-3 days to complete if done from scratch.
Assuming that there are continuous rotations that must be very precisely approximated may not always be necessary. In some cases, as in early demonstrations of FT hardware for example, lower levels of precision (on the order of $\epsilon=10^{-3}$) may be acceptable. In these scenarios, pure search-based tools such as Synthetiq are a better choice. Exhaustively searching the space of circuit implementations may be viable in this case, and it is likely to produce lower gate count results.
Determining exactly what precision is needed to ensure circuits produce meaningful results is problem dependent, and an area that deserves deeper investigation. Analytical decompositions can always be used to produce higher precision approximations, but because these decompositions require so many gates, new techniques must be invented to handle this task. We expect that other analytical-search-based hybrid approaches are possible and practical.
VI-B Interpreting the Value Function
In order for a function to meaningfully guide search-based synthesis, it should correlate with the goal of finding short gate count implementations of unitaries. Intuitively, the value of a unitary state should depend on the number of gates needed to transform that state to a terminal state.
A discrete gate set synthesis algorithm which relies directly on the Hilbert-Schmidt distance function to guide the placement of gates will randomly search the space of possible circuits. Figure 9(c) shows how for random Clifford+T circuits, Hilbert-Schmidt distance to the identity does not correlate with the number of gates in the circuit (and therefore the number of gates needed to synthesize that unitary). Synthesis states in the neighborhood of the current state appear equally distant from a terminal state.
How does a Reinforcement Learning agent avoid this pitfall? Under a reward function such as Equation 7, the learned value of states correlates with the number of gates in the circuit. As shown in Figure 9(d), the value of a state is a learned, scaled and, shifted representation of the “gates to go”. This property makes the agent a fast and efficient synthesizer; it ideally takes the shortest direct path to a terminal state.
In the case of exact synthesis of single-qubit Clifford+T unitaries, the smallest divisor exponent (SDE) plays a similar role as the value function [12]. This synthesis algorithm sequentially takes actions depending on which gate decreases the SDE. In the multi-qubit case, it is believed that such a simple function does not naturally exist [17]. The model-based RL approach taken in this paper can be construed as a learned version of the SDE function for approximate multi-qubit synthesis.
VI-C Scaling to Wider and Deeper Circuits
Only cases of 1-3 qubit synthesis are considered in this paper. Circuits of interest are typically on the order of 100-1000s of logical qubits. As the size the unitaries associated with these circuits grows exponentially in the number of qubits, it is not likely that circuits larger than 10s of qubits wide will ever be directly synthesizable. Past works have addressed this issue with circuit partitioning; by splitting large circuits into smaller subcircuits which can each be individually synthesized [51, 52, 53, 54]. Circuit partitioning is applicable when a circuit implementation of a program is already known. This is a fair assumption to make in the case of gate set transpilation and optimization. Our experiments in Section V have demonstrated that an RL diagonalizing agent is capable of synthesizing high precision unitaries taken directly from partitioned circuits of interest.
In non-discrete synthesis scenarios (i.e. with a get set of parameterized single-qubit unitaries and CNOTs as opposed to Clifford+T), numerical methods have proved invaluable to finding compact circuit implementations of unitaries [55, 56, 57]. Recent work has investigated how these methods can be transferred to the setting of Clifford+T synthesis [58]. As this approach requires solving a complex non-linear constrained optimization problem, its run time can quickly grow unwieldy; synthesis with DVNU requires on the order of seconds as opposed to tens of minutes. As our diagonalization method runs much faster, we foresee that it may be used as a “first pass” optimization method in conjunction with these more powerful synthesizers; cases that cannot be easily diagonalized can instead be handled by constrained numerical optimization or analytical decompositions.
Our method requires only a representation of the state to determine values. This means that the value of placing composite gates can be determined as easily as gates directly represented in the gate set. Scaling to deeper circuits may therefore be possible if a set of common composite gates are collected.
VII Conclusion
We have demonstrated a novel approach to high precision unitary synthesis using discrete gate sets. By diagonalizing unitaries instead of directly inverting them, continuous degrees of freedom can be avoided then handled efficiently using mathematical decomposition methods. This enables us to handle more unitaries than other search-based synthesis methods which directly invert continuous rotations which is intractable for high levels of precision. We developed novel neural architectural components: global phase invariance and unitary perturbations, which enable deep learning entities to more effectively learn on unitary datasets. The effectiveness of our Reinforcement Learning based diagonalizing approach is demonstrated by synthesizing unitaries taken from a suite of partitioned quantum algorithms. Compared to the Quantum Shannon Decomposition, our approach reduces the number of non-trivial rotations (and therefore T-gates) in two-qubit unitaries by $82.8\%$ and $94.8\%$ in three-qubit unitaries. Because the diagonalizing approach is fast and able to find low gate count implementations of meaningful unitaries, it remains a promising technique for use in future fault-tolerant quantum compilers.
VIII Appendix
VIII-A Diagonal Distance
Our goal is to determine when a state $s_{t}$ can be nearly inverted (satisfy the distance condition in Equation 6) by a completely diagonal unitary. Assume the state of the synthesis process is $s_{t}$, and it is “nearly diagonal.” This means that each row of $s_{t}$ is of the form
$s_{t}[i]=\begin{bmatrix}\delta_{i1}&\dots&x_{ii}&\dots&\delta_{i2^{n}}\end{bmatrix}.$
Because $s_{t}$ is unitary, we know that $|x_{ii}|=\sqrt{1-\sum_{j\neq i}\delta_{ij}^{2}}$. Consider a diagonal unitary matrix $D$, and the product $D\times s_{t}$. Rearranging the Hilbert-Schmidt distance function
$\sqrt{4^{n}(1-\epsilon^{2})}\leq|tr(D\times s_{t})|\leq\sum_{i}|d_{ii}|\cdot|x_{ii}|\leq\sum_{i}|x_{ii}|$
then it is the case that
$2^{n}\sqrt{(1-\epsilon^{2})}\leq\sum_{i=1}^{2^{n}}\sqrt{1-\sum_{j\neq i}\delta_{ij}^{2}}\leq\max_{i\in[2^{n}]}2^{n}\sqrt{1-\sum_{j\neq i}\delta_{ij}^{2}}$
which can be reduced to
$\max_{i\in[2^{n}]}\sqrt{\sum_{j\neq i}\delta_{ij}^{2}}\leq\epsilon.$
Thus, if the maximum magnitude of non-diagonal row elements is less than the synthesis threshold, state $s_{t}$ can be multiplied by a purely diagonal matrix so that the Hilbert-Schmidt distance condition is met.
VIII-B Global Phase Invariance
For unitary $U$, we define $\mu=\frac{1}{N^{2}}\sum u_{ij}\in\mathbb{C}$ to be the element-wise mean. As it may be the case that $\mu=0$, we consider the mean of the element-wise squared matrix $U^{\odot 2}$ which is $\mu^{\prime}=\frac{1}{N^{2}}\sum u_{ij}^{2}$. We normalize this to obtain $\hat{\mu^{\prime}}=\mu^{\prime}/\|\mu^{\prime}\|$. Next we halve the phase of $\hat{\mu}^{\prime}$ and conjugate to receive $\hat{\mu}^{*}$. Finally we multiply by $U$ and conditionally add a phase correction of $\pi$ to receive a global phase invariant representation $W$. This process is depicted in Figure 4.
Proof.
Say $U=e^{i\theta}W$. The element-wise mean of $U^{\odot 2}$ is $\mu^{\prime}=e^{i(2\theta+2\pi k)}x$. Normalizing, halving the phase, and conjugating gives $\hat{\mu}^{*}=e^{-i(\theta+k\pi)}$. Multiplying $U$ by the conjugated normalized mean gives $\hat{\mu}^{*}U=e^{i\theta}e^{-i(\theta+k\pi)}W=e^{-ik\pi}W$. From this final result we cannot detect whether $k$ is odd or even. We choose to ensure that the first maximum magnitude entry of the first row of $W$ should have a positive real component. If it does not, we apply the $\pi$ phase correction. ∎